celticcat:

What is the largest value of n for which n! has fewer that 100 decimal digits?

undeadknight26:

I believe the answer is 70, not 100% sure.

sillybilly123:

reckon it's 69, using the Stirling approximation, calculator and trial and error - all of which can then be checked on Wolfram(:-/). is that how you are supposed to be doing this?! i initially thought it was a mod arithmetic thing but can't seem to get that started.

sillybilly123:

\(10^1\) has 2 digits, \(10^2\) has 3, etc ... so we are looking for n where \(\log n! = 99.xxxxxx\) to get a 100 digit factorial Take Stirling: \(n! \approx \left(\dfrac{n}{e} \right)^n\sqrt{2 \pi n}\) re-write it in terms of \(\log\) **base 10**, as: \(\log n! \approx n \log \dfrac{n}{ e} + \log \sqrt{2 \pi n}\) then **trial and error** in calculator gives: - \(\log 69! \approx 98.2\), 99 digits - \(\log 70! \approx 100.1\), ie 101 digits No 100 digit factorial....

celticcat:

Thanks